# Holon metal, charge-density-wave and chiral superconductor from doping fractional Chern insulator and SU(3)1 chiral spin liquid

### v1

@@ III.3 Pairing instability and charge-density-wave trion Fermi liquid
The transverse gauge field then mediates repulsive or attractive interactions between the fermion $\psi_{I}$. We find that $\psi_{1}$ couples to $a_{+}+a_{-}$, $\psi_{2}$ couples to $a_{+}-a_{-}$ and $\psi_{3}$ couples to $-2a_{-}$. It is then easy to find that the intra-flavor BCS interaction $V_{\mathrm{intra}}(\mathbf{q})\sim\frac{4\upsilon_{F}^{2}}{3\chi_{d}q^{2}}$ and the inter-flavor BCS interaction $V_{\mathrm{inter}}(\mathbf{q})\sim-\frac{2\upsilon_{F}^{2}}{3\chi_{d}q^{2}}$. Therefore we find a strong attractive interaction between each pair of Fermi surfaces.
@@
where another internal U(1) gauge field $\gamma_{+}$ is introduced to describe the pairing $\psi_{1}\psi_{2}$, which is charged under the gauge field $-2A_{c}+\alpha_{\beta}$. The integration of $\gamma_{+}$ leads to $\alpha+\beta=2A_{c}$. Then $\psi_{3}$ couples to $-3A_{c}$, consistent with the expectation that $\psi_{3}$ is now a physical charge $-3e$ operator. We are still left with the Chern-Simons terms. We can substitute $\alpha\rightarrow A_{c}+a$ and $\beta\rightarrow A_{c}-a$, then we reach the final theory:
@@
It is interesting to see that this unconventional metallic phase has a Fermi surface formed by charge $-3e$ trion, coexisting with a Hall conductivity $\sigma_{xy}=-6\frac{e^{2}}{h}$. There is also a hidden topological order, which is the same as the SU(2)1 chiral spin liquid. We can also consider other pairing symmetry, but it only influences the hidden topological order and does not alter the transport property.
@@ IV Dope FCI:holon metal and CDW metal
We can now move to the discussion of doping a $\nu=-\frac{2}{3}$ FCI in the context of twisted MoTe2. At $\nu=-1$, we have a Chern insulator with $C=1$. Then at $\nu=-\frac{2}{3}=-1+\frac{1}{3}$, we have an additional $1/3$ Laughlin state from electrons at $1/3$ filling of a $C=-1$ Chern band. This gives a Hall conductivity $\sigma_{xy}=\frac{2}{3}\frac{e^{2}}{h}$ at $\nu=-\frac{2}{3}$. Next we dope the FCI to the filling $\nu=-\frac{2}{3}-x=-1+\frac{1}{3}-x$. It can be understood as doping holes into the $\nu=-\frac{1}{3}$ Laughlin state and thus we expect similar physics as the doped SU(3)1 CSL.
@@
$A$ is the probing field for the electron. One can see the similarity to Eq. 10, but with the mapping $A_{c}=\frac{1}{3}A$. We also replace the background term $\frac{3}{4\pi}A_{c}dA_{c}$ in Eq. 10 with $\frac{1}{4\pi}AdA$, which is from the $\nu=-1$ Chern insulator.
The action above is the standard effective theory of the Laughlin state. Each $\varphi_{I}$ creates an anyon with physical charge $-\frac{e}{3}$. At physical hole density $x$, each $\varphi_{I}$ is at density $n_{\varphi_{I}}=x$. Variation of $a_{0}$ then leads to $\frac{\langle da\rangle}{2\pi}=-\frac{1}{3}\sum_{I=1,2,3}n_{\varphi_{I}}=-x$. So each $\varphi_{I}$ is at effective magnetic filling $\nu_{m}=-1$.
Again we need to deal with boson at $\nu_{m}=-1$. The translation symmetry acts as $T_{x}:\varphi_{I}\rightarrow\varphi_{I}$, $T_{y}:\varphi_{1}\rightarrow\varphi_{2},\varphi_{2}\rightarrow\varphi_{3}$. Following the discussions in doped SU(3) CSL, one possible translation-invariant state is to put each $\varphi_{I}$ in the bosonic CFL phase. We expect a similar holon metal phase as in Eq. 14, but with different charge assignment due to the $A_{c}=3A$ correspondence. Let us label the composite fermion corresponding to each $\varphi_{I}$ as $\psi_{I}$, the low energy theory is:
@@
The phase is similar to the ‘secondary CFL’ in Ref. [40], but there are differences in the coefficient of the self Chern-Simons term. The corresponding phase was dubbed as holon metal in the context of doped SU(3) CSL. Here we still call it holon metal just for simplicity. We note the property of the phase is very different from a CFL. Actually it can also be constructed in a simpler way: we do the part construction of the electron operator $c=f_{1}f_{2}f_{3}$. At $\nu=-1+\frac{1}{3}$, each $f_{I}$ is at density $n_{f}=\frac{1}{3}$. Then we let each $f_{I}$ form a Chern insulator with $C=-1$ with a tripled unit cell in the mean field level. This is the well-known construction of the Laughlin state. Then at hole doped case with $\nu=\frac{2}{3}-x$, each $f_{I}$ is at density $n_{f}=\frac{1}{3}-x$ and can form a small hole pocket with size $x$ on top of the $C=-1$ Chern insulator. The resulting wavefunction is a product of three Slater determinants. We can identify $\psi_{I}$ as the low energy field operator of the hole of $f_{I}$: $\psi_{I}\sim f_{I}^{\dagger}$. One can also derive the same low energy field theory in this new construction (see Appendix. A). Basically we can think in terms of the following simple picture: each hole is split to three fractionalized fermions with physical charge $-\frac{e}{3}$, which form three hole pockets on top of Chern insulators. Naively the mean field ansatz need to triple the unit cell. But the final phase is still translation invariant, just with a projective implementation: $T_{x}:\psi_{I}\rightarrow\psi_{I}$, $T_{y}:\psi_{1}\rightarrow\psi_{2},\psi_{2}\rightarrow\psi_{3},\psi_{3}\rightarrow\psi_{1}$.
@@ IV.3 Stabilization of the holon metal and quantum oscillations under magnetic field
Assuming translation symmetry, we have $n_{1}=n_{2}=n_{3}$ and therefore $b_{\alpha}=b_{\beta}=-\frac{1}{3}B$. Hence there will be spontaneously generated internal flux such that each fermion $\psi_{I}$ feels an effective magnetic field of $-\frac{1}{3}B$, just like a particle with charge $-e/3$.
Because of the effective magnetic field, the inter-flavor pairing $\langle\psi_{1}\psi_{2}\rangle$ should be suppressed when $B$ is larger than a critical value. Therefore the holon metal may be the ground state at finite magnetic field. Then we expect quantum oscillations in the form of $\cos\big{(}\frac{\hbar A_{\mathrm{FS}}}{e}\frac{1}{\frac{1}{3}B}-2\pi\gamma\big{)}$. $\gamma$ is a phase shift from the background Berry phase. Thus the quantum oscillation frequency is $f=\frac{3\hbar A_{\mathrm{FS}}}{e}$. Here $A_{\mathrm{FS}}$ is the Fermi surface volume corresponding to a density $x$ per original unit cell. Note that this frequency is three times larger than a conventional CDW metal with hole pocket formed by charge $-e$ hole.
@@ V Discussion on the experiment in twisted MoTe2
From the above analysis, we can see that the most natural symmetric non-quantum-Hall state upon doping the $\nu=-\frac{2}{3}$ FCI is a holon metal, which is then unstable to a CDW metal at lower temperature. On the other hand, we do not find a natural way to enter an ‘anyon superconductor’ phase directly from an anyon gas, if the cheapest anyon is the elementary $1/3$ anyon. Nevertheless, signatures of superconductivity were reported from doping the $\nu=-\frac{2}{3}$ FCI in twisted MoTe2[42]. Here we briefly discuss possible scenarios, without touching the microscopic pairing mechanism.
@@
(III) In the more exotic scenario, one may imagine that the cheapest excitation of the FCI is the $2/3$ anyon for some unknown reasons. Although this is quite counterintuitive, there exist some numerical evidences for fractional quantum Hall state[49]. Ref.[40] showed that a superconductor equivalent to four copies of $p-ip$ superconductor can emerge in this scenario. In Appendix. D, we also explore this scenario in our framework and find a slightly more exotic superconductor. It also has chiral central charge of $c=-2$, but hosts a deconfined anti-semion. In this exotic scenario, the normal state of the superconductor is very unconventional and is essentially a gas of $2/3$ anyons.
Figure 1: (a) Illustration of phase diagram with temperature and doping $n_{e}=\frac{1}{3}-x$ at small $x$. ‘SC’ indicates superconductor and ‘CDW’ indicates the CDW metal. Anyon gas can just crossover to the holon metal without any sharp phase transition. The holon metal is unstable to the CDW metal. A superconductor, if exists, may likely emerge within the CDW metal phase, but the microscopic pairing mechanism is unclear. (b) At a fixed $x$, an illustrated phase diagram with the temperature $T$ and the magnetic field $B$. In the holon metal phase, each pocket from $\psi_{I}$ feels an effective magnetic field of $-\frac{1}{3}B$, which can suppress the pairing and the instability into the CDW phase.
In principle there can be a more trivial scenario: the superconductor is from a conventional Fermi liquid with an electron pocket of size $1/3-x$ and no symmetry breaking, as discussed in the weak coupling theory[50]. In this scenario, it is not clear how the Fermi liquid evolves to the FCI or anyon gas phase without an intermediate step. So far, the only continuous evolution we know from FL to FCI is through a CDW phase[28]. Therefore a large Fermi surface state at small doping away from the FCI phase may not be very natural, unless there is a first order transition. We leave it to the future experiments to measure the Hall number in the normal state, which can easily distinguish small hole pocket in our holon metal or CDW metal versus the large electron pocket in the conventional Fermi liquid.
@@ VI Conclusion
In summary, we provide a unified theoretical description for doped fractional Chern insulator at $\nu=-\frac{2}{3}$ and doped SU(3)1 chiral spin liquid. Assuming a finite spin gap, doped SU(3)1 is dual to the doped FCI with a spin-singlet charge $3e$ trion corresponding to the electron in the FCI case. In both cases, a natural possibility is a holon metal phase with three small Fermi surfaces. Each pocket is formed by a spinless charge $-e$ holon in the doped CSL case and a $-e/3$ holon in the doped FCI case. In both cases the holon metal is likely unstable to a CDW metal, but it may be stabilized by externally applied magnetic field. A superconductor, as observed in the twisted MoTe2 system[42], may emerge from the holon metal or CDW metal as its normal state. These normal states have small hole Fermi surfaces and are thus very different from the weak coupling theory assuming a large electron Fermi surface.
Acknowledgement I thank Zhengyan Darius Shi for pointing out Ref.49 and the possibility of cheaper charge $32e/$ 3 anyon. I thank Xue-Yang Song and Ashvin Vishwanath on earlier collaborations on doping chiral spin liquid. The work is supported by the National Science Foundation under Grant No. DMR-2237031.
@@ Appendix A Transport property for holon metal from doping the FCI
Let’s provide an alternative derivation of the effective theory of the holon metal from hole doping the FCI at $\nu=-\frac{2}{3}$. On top of the $\nu=-1$ Chern insulator, we have electron at density $n_{e}=\frac{1}{3}$. We then do the parton construction of the electron operator: $c(\mathbf{r})=f_{1}(\mathbf{r})f_{2}(\mathbf{r})f_{3}(\mathbf{r})$. This gives a SU(3) gauge redundancy, but we assume our mean field ansatz higgses it down to $U(1)\times U(1)$. There are two U(1) gauge fields $\alpha_{\mu}$ and $\beta_{\mu}$. We also introduce the probing field $A_{\mu}$ and assign each $f_{I}$ to have charge $1/3$ under $A_{\mu}$. Note this charge assignment can be shifted by redefining $\alpha$ and $\beta$. As discussed in the main text, we can put each $f_{I}$ in a mean field ansatz with a tripled unit cell such that it can form a Chern insulator with $C=-1$ at filling $n_{f}=\frac{1}{3}$. Now at filling $n_{f}=\frac{1}{3}-x$, each $f_{I}$ just forms an additional hole pocket. Let us first assume that each $f_{I}$ has an ansatz with uniform flux of $\frac{2\pi}{3}$ per triangular unit cell, so that there are three equivalent minima under hole doping. We label the hole-creation operator of these three minima as $\psi_{I;a}=f_{I}^{\dagger}$ with $I=1,2,3$ and $a=1,2,3$. Here $a$ is the valley index that labels the three minima. In principle each $f_{I}$ can form three hole pockets, occupying all of the three valleys. However, we can reach a simpler ansatz by adding perturbations which split the energy of these three valleys so that $f_{I}$ forms a single hole pocket corresponding to $a=I$. Naively, this breaks the translation symmetry. However, we can always combine an exchange between the three $f_{I}$ in translation symmetry. More precisely, we have the following implementation of the translation symmetry: $T_{x}:\psi_{I;a}\rightarrow\psi_{I;a}$ and $T_{y}:\psi_{I;a}\rightarrow\psi_{I+1;a+1}$. Here $I+1$ and $a+1$ are defined as mod $3$ so $3+1=1$. One can check that a perturbation term like $\delta H=-\Delta\sum_{I=1,2,3}\psi^{\dagger}_{I;I}\psi_{I;I}$ is allowed. Then in the low energy we have one single hole pocket from each $f_{I}$, formed by the low energy field operator $\psi_{I}=\psi_{I;I}$.
@@
Next we derive the low energy effective theory. In our theory $f_{1}$ couples to $\frac{1}{3}A_{\mu}-\alpha_{\mu}$, $f_{2}$ couples to $\frac{1}{3}A_{\mu}-\beta_{\mu}$ and $f_{3}$ couples to $\frac{1}{3}A_{\mu}+\alpha_{\mu}+\beta_{\mu}$. $c=f_{1}f_{2}f_{3}$ is gauge invariant and only couples to $A_{\mu}$. Because each $f_{I}$ has a background Chern insulator, we get a background Chern-Simons term: $-\frac{1}{4\pi}(\frac{1}{3}A-\alpha)d(\frac{1}{3}A-\alpha)-\frac{1}{4\pi}(\frac{1}{3}A-\beta)d(\frac{1}{3}A-\beta)-\frac{1}{4\pi}(\frac{1}{3}A+\alpha+\beta)d(\frac{1}{3}A+\alpha+\beta)$. The final theory is:

### v2

@@ III.3 Pairing instability and charge-density-wave trion Fermi liquid
The transverse gauge field then mediates repulsive or attractive interactions between the fermion $\psi_{I}$. We find that $\psi_{1}$ couples to $a_{+}+a_{-}$, $\psi_{2}$ couples to $a_{+}-a_{-}$ and $\psi_{3}$ couples to $2a_{-}$. It is then easy to find that the intra-flavor BCS interaction $V_{\mathrm{intra}}(\mathbf{q})\sim\frac{4\upsilon_{F}^{2}}{3\chi_{d}q^{2}}$ and the inter-flavor BCS interaction $V_{\mathrm{inter}}(\mathbf{q})\sim-\frac{2\upsilon_{F}^{2}}{3\chi_{d}q^{2}}$. Therefore we find a strong attractive interaction between each pair of Fermi surfaces.
@@
where another internal U(1) gauge field $\gamma_{+}$ is introduced to describe the pairing $\psi_{1}\psi_{2}$, which is charged under the gauge field $2A_{c}+\alpha_{\beta}$. The integration of $\gamma_{+}$ leads to $\alpha+\beta=2A_{c}$. Then $\psi_{3}$ couples to $3A_{c}$, consistent with the expectation that $\psi_{3}$ is now a physical charge 3e-3e operator. We are still left with the Chern-Simons terms. We can substitute $\alpha\rightarrow A_{c}+a$ and $\beta\rightarrow A_{c}-a$, then we reach the final theory:
@@
It is interesting to see that this unconventional metallic phase has a Fermi surface formed by charge 3e-3e trion, coexisting with a Hall conductivity $\sigma_{xy}=-6\frac{e^{2}}{h}$. There is also a hidden topological order, which is the same as the SU(2)1 chiral spin liquid. We can also consider other pairing symmetry, but it only influences the hidden topological order and does not alter the transport property.
@@ IV Dope FCI:holon metal and CDW metal
We can now move to the discussion of doping a $\nu=-\frac{2}{3}$ FCI in the context of twisted MoTe2. At $\nu=-1$, we have a Chern insulator with $C=1$. Then at $\nu=-\frac{2}{3}=-1+\frac{1}{3}$, we have an additional $1/3$ Laughlin state from electrons at $1/3$ filling of a C=1C=-1 Chern band. This gives a Hall conductivity $\sigma_{xy}=\frac{2}{3}\frac{e^{2}}{h}$ at $\nu=-\frac{2}{3}$. Next we dope the FCI to the filling $\nu=-\frac{2}{3}-x=-1+\frac{1}{3}-x$. It can be understood as doping holes into the $\nu=-\frac{1}{3}$ Laughlin state and thus we expect similar physics as the doped SU(3)1 CSL.
@@
AA is the probing field for the electron. One can see the similarity to Eq. 10, but with the mapping $A_{c}=\frac{1}{3}A$. We also replace the background term $\frac{3}{4\pi}A_{c}dA_{c}$ in Eq. 10 with $\frac{1}{4\pi}AdA$, which is from the $\nu=-1$ Chern insulator.
The action above is the standard effective theory of the Laughlin state. Each $\varphi_{I}$ creates an anyon with physical charge $\frac{e}{3}$. At physical hole density $x$, each $\varphi_{I}$ is at density $n_{\varphi_{I}}=x$. Variation of $a_{0}$ then leads to $\frac{\langle da\rangle}{2\pi}=-\frac{1}{3}\sum_{I=1,2,3}n_{\varphi_{I}}=-x$. So each $\varphi_{I}$ is at effective magnetic filling $\nu_{m}=-1$.
Again we need to deal with boson at $\nu_{m}=-1$. The translation symmetry acts as $T_{x}:\varphi_{I}\rightarrow e^{i\frac{2\pi}{3}I}\varphi_{I}$, $T_{y}:\varphi_{1}\rightarrow\varphi_{2},\varphi_{2}\rightarrow\varphi_{3}$. Following the discussions in doped SU(3) CSL, one possible translation-invariant state is to put each $\varphi_{I}$ in the bosonic CFL phase. We expect a similar holon metal phase as in Eq. 14, but with different charge assignment due to the $A_{c}=3A$ correspondence. Let us label the composite fermion corresponding to each $\varphi_{I}$ as $\psi_{I}$, the low energy theory is:
@@
The phase is similar to the ‘secondary CFL’ in Ref. [40], but there are differences in the coefficient of the self Chern-Simons term. The corresponding phase was dubbed as holon metal in the context of doped SU(3) CSL. Here we still call it holon metal just for simplicity. We note the property of the phase is very different from a CFL. Actually it can also be constructed in a simpler way: we do the parton construction of the electron operator $c=f_{1}f_{2}f_{3}$. At $\nu=-1+\frac{1}{3}$, each $f_{I}$ is at density $n_{f}=\frac{1}{3}$. Then we let each $f_{I}$ form a Chern insulator with C=1C=-1 with a tripled unit cell in the mean field level. This is the well-known construction of the Laughlin state. Then at hole doped case with $\nu=\frac{2}{3}-x$, each $f_{I}$ is at density $n_{f}=\frac{1}{3}-x$ and can form a small hole pocket with size $x$ on top of the C=1C=-1 Chern insulator. The resulting wavefunction is a product of three Slater determinants. We can identify $\psi_{I}$ as the low energy field operator of the hole of $f_{I}$: $\psi_{I}\sim f_{I}^{\dagger}$. One can also derive the same low energy field theory in this new construction (see Appendix. A). Basically we can think in terms of the following simple picture: each hole is split to three fractionalized fermions with physical charge $\frac{e}{3}$, which form three hole pockets on top of Chern insulators. Naively the mean field ansatz need to triple the unit cell. But the final phase is still translation invariant, just with a projective implementation: $T_{x}:\psi_{I}\rightarrow e^{i\frac{2\pi}{3}I}\psi_{I}$, $T_{y}:\psi_{1}\rightarrow\psi_{2},\psi_{2}\rightarrow\psi_{3},\psi_{3}\rightarrow\psi_{1}$.
@@ IV.3 Stabilization of the holon metal and quantum oscillations under magnetic field
Assuming translation symmetry, we have $n_{1}=n_{2}=n_{3}$ and therefore $b_{\alpha}=b_{\beta}=-\frac{1}{3}B$. Hence there will be spontaneously generated internal flux such that each fermion $\psi_{I}$ feels an effective magnetic field of $\frac{1}{3}B$, just like a particle with charge e/3-e/3.
Because of the effective magnetic field, the inter-flavor pairing $\langle\psi_{1}\psi_{2}\rangle$ should be suppressed when $B$ is larger than a critical value. Therefore the holon metal may be the ground state at finite magnetic field. Then we expect quantum oscillations in the form of $\cos\big(\frac{\hbar A_{\mathrm{FS}}}{e}\frac{1}{\frac{1}{3}B}-2\pi\gamma\big)$. $\gamma$ is a phase shift from the background Berry phase. Thus the quantum oscillation frequency is $f=\frac{3\hbar A_{\mathrm{FS}}}{e}$. Here $A_{\mathrm{FS}}$ is the Fermi surface volume corresponding to a density $x$ per original unit cell. Note that this frequency is three times larger than a conventional CDW metal with hole pocket formed by charge e-e hole.
@@ V Discussion on the experiment in twisted MoTe2
In our framework, the most natural symmetric non-quantum-Hall state upon doping the $\nu=-\frac{2}{3}$ FCI is a holon metal, which is then unstable to a CDW metal at lower temperature. On the other hand, we do not find a natural way to enter an ‘anyon superconductor’ phase directly from an anyon gas, if the cheapest anyon is the elementary one with charge $e/3$. Nevertheless, signatures of superconductivity were reported from doping the $\nu=-\frac{2}{3}$ FCI in twisted MoTe2[42]. Here we briefly discuss possible scenarios, without touching the microscopic pairing mechanism.
@@
(III) In the more exotic scenario, one may imagine that the cheapest excitation of the FCI is the $2/3$ anyon for some unknown reasons. Although this is quite counterintuitive, there exist some numerical evidences to support it in fractional quantum Hall state[49]. Ref.[40] showed that a superconductor equivalent to four copies of pipp-ip superconductor can emerge in this scenario. In Appendix. D, we also explore this scenario in our framework and find a slightly more exotic superconductor. It also has chiral central charge of c=2c=-2, but hosts a deconfined anti-semion. In this exotic scenario, the normal state of the superconductor is very unconventional and is essentially a gas of $2/3$ anyons.
Figure 1: (a) Illustration of phase diagram with temperature and doping $n_{e}=\frac{1}{3}-x$ at small $x$. ‘SC’ indicates superconductor and ‘CDW’ indicates the CDW metal. Anyon gas can just crossover to the holon metal without any sharp phase transition. The holon metal is unstable to the CDW metal. A superconductor, if exists, may likely emerge within the CDW metal phase, but the microscopic pairing mechanism is unclear. (b) At a fixed $x$, an illustrated phase diagram with the temperature $T$ and the magnetic field $B$. In the holon metal phase, each pocket from $\psi_{I}$ feels an effective magnetic field of $\frac{1}{3}B$, which can suppress the pairing and the instability into the CDW phase.
(IV) In the more conventional scenario, the superconductor is from a conventional Fermi liquid with an electron pocket of size 1/3x1/3-x and no symmetry breaking, as discussed in the weak coupling theory[50, 51]. In this scenario, it is not clear how the Fermi liquid evolves to the FCI or anyon gas phase without an intermediate step. It is possible that the superconductivity instability is enhanced in a strongly correlated Fermi liquid proximate to the FCI phase[52]. Ref. [53] suggests that Fermi liquid phase close to the FCI may have an additional factor $|z_{i}-z_{j}|^{6}$ in its wavefunction. It is interesting to explore whether an effective attractive interaction arises for the quasi-particles in such a case.
We leave it to the future experiments to measure the Hall number in the normal state, which can easily distinguish small hole pocket in our holon metal or CDW metal versus the large electron pocket in the conventional Fermi liquid.
@@ VI Conclusion
In summary, we provide a unified theoretical description of the doped fractional Chern insulator (FCI) at $\nu=-\frac{2}{3}$ and the doped SU(3)1 chiral spin liquid. Assuming a finite spin gap, the doped SU(3)1 system is dual to the doped FCI, where a spin-singlet charge-$3e$ trion corresponds to the electron in the FCI limit. In both regimes, a natural candidate phase is a holon metal featuring three small Fermi pockets. These pockets are formed by spinless holons carrying charge e-e in the doped CSL and charge e/3-e/3 in the doped FCI. While this holon metal is likely unstable toward a charge density wave (CDW) state, it may be stabilized by an external magnetic field. We leave the investigation of the interplay between the holon metal and potential superconductivity to future work
Acknowledgement I thank Zhengyan Darius Shi for pointing out the possibility of cheap charge $2e/3$ anyon. I thank Xue-Yang Song and Ashvin Vishwanath for earlier collaborations on doping SU(2) chiral spin liquid. The work is supported by the National Science Foundation under Grant No. DMR-2237031.
@@ Appendix A Transport property for holon metal from doping the FCI
Let’s provide an alternative derivation of the effective theory of the holon metal from hole doping the FCI at $\nu=-\frac{2}{3}$. On top of the $\nu=-1$ Chern insulator, we have electron at density $n_{e}=\frac{1}{3}$. We then do the parton construction of the electron operator: $c(\mathbf{r})=f_{1}(\mathbf{r})f_{2}(\mathbf{r})f_{3}(\mathbf{r})$. This gives a SU(3) gauge redundancy, but we assume our mean field ansatz higgses it down to $U(1)\times U(1)$. There are two U(1) gauge fields $\alpha_{\mu}$ and $\beta_{\mu}$. We also introduce the probing field $A_{\mu}$ and assign each $f_{I}$ to have charge $1/3$ under $A_{\mu}$. Note this charge assignment can be shifted by redefining $\alpha$ and $\beta$. As discussed in the main text, we can put each $f_{I}$ in a mean field ansatz with a tripled unit cell such that it can form a Chern insulator with C=1C=-1 at filling $n_{f}=\frac{1}{3}$. Now at filling $n_{f}=\frac{1}{3}-x$, each $f_{I}$ just forms an additional hole pocket. Let us first assume that each $f_{I}$ has an ansatz with uniform flux of $\frac{2\pi}{3}$ per triangular unit cell, so that there are three equivalent minima under hole doping. We label the hole-creation operator of these three minima as $\psi_{I;a}=f_{I}^{\dagger}$ with $I=1,2,3$ and $a=1,2,3$. Here $a$ is the valley index that labels the three minima. In principle, each $f_{I}$ can form three hole pockets, occupying all three valleys. However, we can reach a simpler ansatz by adding perturbations which split the energy of these three valleys so that $f_{I}$ forms a single hole pocket corresponding to $a=I$. Naively, this breaks the translation symmetry. However, we can always combine an exchange between the three $f_{I}$ in translation symmetry. More precisely, we have the following implementation of the translation symmetry: $T_{x}:\psi_{I;a}\rightarrow e^{i\frac{2\pi}{3}I}\psi_{I;a}$ and $T_{y}:\psi_{I;a}\rightarrow\psi_{I+1;a+1}$. Here $I+1$ and $a+1$ are defined as mod $3$ so $3+1=1$. One can check that a perturbation term like $\delta H=-\Delta\sum_{I=1,2,3}\psi^{\dagger}_{I;I}\psi_{I;I}$ is allowed. Then in the low energy we have one single hole pocket from each $f_{I}$, formed by the low energy field operator $\psi_{I}=\psi_{I;I}$.
@@
Next we derive the low energy effective theory. In our theory $f_{1}$ couples to $\frac{1}{3}A_{\mu}-\alpha_{\mu}$, $f_{2}$ couples to $\frac{1}{3}A_{\mu}-\beta_{\mu}$ and $f_{3}$ couples to $\frac{1}{3}A_{\mu}+\alpha_{\mu}+\beta_{\mu}$. $c=f_{1}f_{2}f_{3}$ is gauge invariant and only couples to $A_{\mu}$. Because each $f_{I}$ has a background Chern insulator, we get a background Chern-Simons term: $\frac{1}{4\pi}(\frac{1}{3}A-\alpha)d(\frac{1}{3}A-\alpha)-\frac{1}{4\pi}(\frac{1}{3}A-\beta)d(\frac{1}{3}A-\beta)-\frac{1}{4\pi}(\frac{1}{3}A+\alpha+\beta)d(\frac{1}{3}A+\alpha+\beta)$. The final theory is: